Let A be the set of even numbers on a twelve-sided dice
and B the set of numbers greater than 4 on a twelve-sided dice.
Which statement is NOT true?

N(B) = 8
N (A U B) = 10
N (A n B) = 2
N(A) = 6

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3838)


Hints
A = {2,4,6,8,10,12}
B = {5,6,7,8,9,10,11,12}
Notation like n(A) means "the number of values in set A".

is the intersection of sets A and B
It's the set of elements found in both sets
For example, if A = {p,q,r} and B = {r,s,t,p} then

Answer by ikleyn(53937)

.
Let A be the set of even numbers on a twelve-sided dice
and B the set of numbers greater than 4 on a twelve-sided dice.
Which statement is NOT true?

(a) N(B) = 8
(b) N (A U B) = 10
(c) N (A n B) = 2
(d) N(A) = 6
~~~~~~~~~~~~~~~~~~~~~~

To answer the problem's question, you should check every option.


(a)  To check if (a) is TRUE, calculate the number of elements in set B.


(b)  To check if (b) is TRUE, calculate the number of elements in set (A U B).


(c)  To check if (c) is TRUE, calculate the number of elements in set (A n B).


(d)  To check if (d) is TRUE, calculate the number of elements in set A.


Each of listed exercises is extremely simple. Use your fingers as your helpers.

------------------

You got precise instructions from me on  " what to do "  and  " how to do ".

So,  boldly go forward.  If on your way forward you will have questions,  let me know.

You may report me about your progress.